### v2

# Make an Offer They Can’t Refuse: Grounding Bayesian Persuasion in Real-World Dialogues without Pre-Commitment

Abstract.
Persuasion, a fundamental social capability for humans, remains a challenge for AI systems such as large language models (LLMs). Current studies often overlook the strategic use of information asymmetry in message design or rely on strong assumptions regarding pre-commitment. In this work, we explore the application of Bayesian Persuasion (BP) in natural language within single-turn dialogue settings, to enhance the strategic persuasion capabilities of LLMs. Our framework incorporates a commitment-communication mechanism, where the persuader explicitly outlines an information schema by narrating their potential types (e.g., honest or dishonest), thereby guiding the persuadee in performing the intended Bayesian belief update. We evaluate two variants of our approach: Semi-Formal-Natural-Language (SFNL) BP and Fully-Natural-Language (FNL) BP, benchmarking them against both naive and strong non-BP (NBP) baselines within a comprehensive evaluation framework. This framework covers a diverse set of persuadees—including LLM instances with varying prompts and fine-tuning and human participants—across tasks ranging from specially designed persuasion scenarios to general everyday situations. Experimental results on LLM-based agents reveal three main findings: (1) LLMs guided by BP strategies consistently achieve higher persuasion success rates than NBP baselines; (2) SFNL exhibits greater credibility and logical coherence, while FNL shows stronger emotional resonance and robustness in naturalistic conversations; (3) with supervised fine-tuning, smaller models can attain BP performance comparable to that of larger models.
Key words and phrases:
Bayesian Persuasion, Information Design, Conversational LLM
1. Introduction
Figure 1. Illustration of Type-Induced NL Bayesian Persuasion. The upper pandel shows the classic BP components, which correspond to the $m_{basic}$, $m_{type}$, $m_{des}$ and $m_{inf}$ in the following two panels. The middle panel shows an explicit verbalization of the sender’s probabilistic schema and expected utility, where the persuader computes the expected payoff using Bayes’ rule. The lower panel translates the same reasoning into more natural conversational language, expressing uncertainty, confidence, and cost–benefit judgment.
[…]
Two dialogue panels comparing mathematical and natural verbalizations of Bayesian Persuasion. The upper dialogue explicitly states conditional probabilities and expected utility calculations. The lower dialogue expresses equivalent reasoning through natural language, showing how LLMs can transform formal Bayesian reasoning into intuitive persuasive communication.
Persuasion is a fundamental form of human social interaction, enabling individuals to influence others’ beliefs and decisions through communication Brembeck (1976). While large language models (LLMs) already exhibit strong abilities in language generation and understanding, they remain limited in strategic persuasion tasks: LLMs often fail to effectively exploit information asymmetry and struggle to design messages that rationally shift a persuadee’s beliefs.
Bayesian Persuasion (BP), a game-theoretic framework for information design, offers a complete solution in constrained mathematical settings by modeling how a persuader can disclose information to maximize desired actions Kamenica and Gentzkow (2011). However, applying BP in open-ended natural language dialogue raises a key challenge: the mathematical constructs of BP—such as priors, world states, and posterior updates—must be verbalized into coherent and persuasive arguments. For example, when convincing friends to try a new restaurant, the priors (the probability that the restaurant will suit their taste) and utilities (how much they would enjoy it) are rarely stated explicitly. The offer would be refused if the story is not that convincing. Moreover, a pivotal step in BP is the sender’s commitment to a signaling schema, which we argue is crucial for natural language realization.
Existing approaches, such as the method in Li et al. (2025), typically rely on pre-commitment by statically encoding the schema in the persuadee’s prompt, bypassing the need for its communication. Consequently, there is a lack of systematic methods for implementing BP where the schema itself is articulated within a single-turn natural language interaction under information asymmetry.
This limitation stems from the inflexibility of pre-commitment in a dynamic dialogue. To overcome this, we introduce a type-induced commitment-communication mechanism: rather than assuming pre-commitment, the persuader explicitly narrates their potential types (e.g., honest or dishonest) within the natural language exchange itself. This verbal articulation of the information schema enables the persuadee to perform Bayesian posterior updates directly from the conversation flow. Thus, we recast the schema pre-commitment as type disclosure, bridging the gap between formal BP theory and authentic natural language implementation within single-turn dialogues.
Motivated by preliminary observations that powerful LLMs such as GPT-5 OpenAI (2025) and DeepSeek-V3.1 DeepSeek-AI (2025a) display BP-like reasoning capability when prompted with above mechanism carefully, we investigate two key questions: (1) Can LLMs consistently execute BP strategies through natural language to enhance persuasion success? And (2) can supervised fine-tuning empower smaller models to achieve BP performance comparable to larger counterparts?
We implement the commitment-communication mechanism in two phases. First, in the explicit view, we provide LLM persuaders with the scenario-specific Bayesian setup (priors, states, utilities) externally, enabling two natural language variants: Semi-Formal-Natural-Language (SFNL) BP, which delivers persuasion through blended calculations and narratives, and Fully-Natural-Language (FNL) BP, which relies solely on fluent discourse without formal computations. Second, in the self-derived view, models infer the Bayesian setup independently, advancing to self-derived SFNL and self-derived FNL to mimic real-world ambiguity. This phased approach allows us to conduct fine-grained analysis and systematically evaluate how BP strategies translate from structured to naturalistic settings.
Our comprehensive evaluation spans diverse persuadees—including LLM instances with varying prompts and fine-tuning, as well as human participants—across tailored persuasion scenarios and everyday contexts. Results demonstrate that: (1) BP-guided LLMs significantly outperform non-BP (NBP) and other strategic baselines in persuasion success, (2) SFNL excels in credibility and logical coherence, whereas FNL induces stronger emotional resonance and robustness, and (3) with supervised fine-tuning, smaller models match the BP performance of larger models.
In summary, this work contributes a framework for implementing BP in natural language dialogues, leveraging explicit commitment to overcome verbalization challenges. By systematically evaluating BP variants, we provide insights into how LLMs can harness information asymmetry for strategic persuasion, paving the way for more effective AI communicators.
2. Related work
2.1. Persuasion with Large Language Models
Recent research has increasingly explored the persuasive capabilities of LLMs, examining both their potential benefits and societal risks. Ramani et al. (2024) investigate multi-agent frameworks that could enhance persuasion efficacy through collaborative specialization, where auxiliary agents handle tasks like strategy development and resistance analysis. Gemp et al. (2024) introduce equilibrium solvers that guide LLM dialogue generation, enabling models to reason about consistent strategic interactions. Shi (2025) reviews AI persuasion systems from a social good perspective, offering theoretical support for discussing ethical implications. Cheng and You (2025) develop theory-driven evaluation frameworks to systematically measure LLMs’ ability to change beliefs and decisions.
Additionally, several studies are focused on building specialized datasets. Persuasion for Good Wang et al. (2020) examines how individual traits affect persuasion outcomes and provides a basis for strategy adaptation. Jin et al. (2024); Hayati et al. (2020) build datasets for the persuasion in daily scenarios. CToMPersu Zhang and Zhou (2025) proposes a “double-blind” framework where persuasion strategies and mental states remain undisclosed, which is particularly well-suited for BP analysis due to its explicit information asymmetry design.
2.2. Bayesian Persuasion and Information Design
As an information design method, BP provides a normative framework for understanding how an informed sender can design signals to influence a receiver’s beliefs and actions. The seminal work of Kamenica and Gentzkow (2011) characterizes when persuasive signaling benefits the sender and derives optimal signaling schemes. This foundation has been expanded through the broader lens of information design Bergemann and Morris (2016, 2019), which unifies BP with communication in games and robust prediction.
Algorithmic perspectives on information design are comprehensively reviewed by Dughmi (2017). Some studies adapt BP to dynamic and uncertain environments. Castiglioni et al. (2020); Bernasconi et al. (2023); Bacchiocchi et al. (2024a) address online extensions of BP and continuously improve the regret bound. Bacchiocchi et al. (2024b) tackle the fully prior-free case. Other directions include persuasion with externalities where one receiver’s actions affect others’ utilities Shaki et al. (2025) and complexity analyses showing that discovering persuasive messages is NP-hard while verifying them is tractable Wojtowicz (2024).
Beyond theoretical extensions, BP has been applied to practical AI challenges. Bai et al. (2024) employ BP for model-agnostic alignment, using a smaller model as an advisor that sends signals to guide larger models’ responses, demonstrating BP’s utility in improving model performance without direct training. Zhang et al. (2025) introduce BP into AI alignment and governance, proposing that AI can be made to act according to human intentions by designing information during the post-deployment phase.
2.3. Verbalizing Bayesian Persuasion in Natural Language
Translating BP into natural language requires verbalizing signals, beliefs, and posterior updates within coherent narratives. Li et al. (2025) address this through Verbalized Bayesian Persuasion (VBP), which incorporates the signaling scheme directly into the receiver’s prompt and adjusts sender generation via keyword manipulation. While effective in their evaluated domains, this approach relies on explicit prompt engineering and is demonstrated in limited, scenario-specific settings.
In contrast, our method enables fully self-derived schema communication: the persuader explicitly articulates the information structure within the natural language discourse itself. This commitment-communication mechanism allows our approach to be model-agnostic, benefiting both large and small models, and generalizable across diverse scenarios using a unified prompting strategy. By operating in single-turn dialogues and supporting both explicit and self-derived Bayesian reasoning, our framework offers broader applicability while maintaining alignment with BP principles.
3. Type-Induced Bayesian Persuasion in Natural Language
In this section, we formalize our approach to one-turn, two-agent persuasion scenarios. Unlike classic BP, our framework addresses more realistic settings where no pre-committed common knowledge exists between agents. Instead, we leverage the richness of natural language to enable the Sender to dynamically frame the interaction beyond classical BP constraints. Crucially, we model a one-turn rhetorical tactic that employs BP logic and language for persuasion, rather than a repeated game with verifiable interactions.
3.1. General Framework
Our framework comprises the following components:
  • Players: A Sender (S) and a Receiver (R).
  • State of the World: A finite set of possible states $\Omega=\{\omega_{1},\dots,\omega_{K}\}$. The true state $\omega\in\Omega$ is observable to S but not to R.
  • Prior Beliefs: R and S hold prior beliefs $\mu_{R},\mu_{S}\in\Delta(\Omega)$ about the world state, respectively. We do not assume these priors are identical or constitute common knowledge.
  • Actions and Utilities: The Sender sends a natural language message $m$. The Receiver observes $m$ and chooses an action $a\in\{\text{Accept, Reject}\}$. The Receiver aims to maximize their expected utility $u(a,\omega)$, while the Sender aims to maximize the probability of acceptance, $P(a=\text{Accept}\mid m)$.
3.2. The Type-Induced Signal
In contrast to the classic BP assumption of a pre-committed, common-knowledge information schema $\pi_{S}(m\mid\omega):\Omega\to\Delta(M)$, where $M$ is a limited set of signal categories, our approach enables dynamic schema conveyance within natural language through type disclosure.
Composite Signal Structure:
The Sender’s message $m$ integrates four functional components:
  • $m_{\text{basic}}$: Background information about the world states ($\Omega$), to align the players’ understanding.
  • $m_{\text{type}}$: A narrative about the Sender’s type, used to construct the information schema.
  • $m_{\text{des}}$: A description of the observed state $\omega$.
  • $m_{\text{inf}}$: An explicit inference, guiding the Receiver to calculate their expected payoff and conclude that accepting is the optimal action.
Type-Induced Information Schema:
The schema emerges from the Sender’s type narrative rather than being pre-defined.
  • Sender Types ($\Theta$): We consider a set of sender types $\Theta=\{\theta_{H},\theta_{D}\}$, representing Honest and Dishonest respectively.
  • Base Policies: Each type is associated with a base communication policy $\pi_{\theta}(m_{\text{des}}\mid\omega)$:
    • The Honest type’s policy, $\pi_{H}(m_{\text{des}}\mid\omega)$, is to truthfully reveal the state $\omega$.
    • The Dishonest type’s policy, $\pi_{D}(m_{\text{des}}\mid\omega)$, is to send a message that maximizes the chance of persuasion, even if it misrepresents the state $\omega$.
  • Schema Induction: The Sender’s utterance $m_{\text{type}}$ induces belief distribution $p(\theta)\in\Delta(\Theta)$ in the Receiver’s mind. For example:
    ”If the car is bad ($\omega_{\text{bad}}$), assume I’m a liar ($\theta_{D}$) 80% of the time, but there’s a 20% chance I’m being honest ($\theta_{H}$).”
    This narrative induces the belief $p(\theta_{D})=0.8$ and $p(\theta_{H})=0.2$. This, in turn, allows the Receiver to infer an effective information schema $\bar{\pi}$ as the weighted average of the base policies:
    $\bar{\pi}(m_{\text{des}}\mid\omega)=p(\theta_{H})\pi_{H}(m_{\text{des}}\mid\omega)+p(\theta_{D})\pi_{D}(m_{\text{des}}\mid\omega)$
3.3. Receiver’s Inference and Decision
After the schema is conveyed via $m_{\text{type}}$, the Receiver uses it to interpret the descriptive signal $m_{\text{des}}$ through the following process:
  • The Receiver observes $m_{\text{des}}$ and uses the effective schema $\bar{\pi}$ to update the prior belief $\mu_{R}$ to a posterior belief $\mu^{\prime}_{R}$.
    $\mu^{\prime}_{R}(\omega)=\frac{\bar{\pi}(m_{\text{des}}\mid\omega)\mu_{R}(\omega)}{\sum_{\omega^{\prime}\in\Omega}\bar{\pi}(m_{\text{des}}\mid\omega^{\prime})\mu_{R}(\omega^{\prime})}$
  • The Receiver selects optimal action $a^{*}$ maximizing expected utility under $\mu^{\prime}_{R}$:
    $a^{*}=\arg\max_{a\in A}\mathbb{E}_{\omega\sim\mu^{\prime}_{R}}[u(a,\omega)]$
  • This entire inference process can be explicitly guided or even performed for the Receiver by the Sender’s utterance $m_{\text{inf}}$.
3.4. Verbalizing the Composite Signal
We implement two verbalization approaches: Semi-Formal-Natural-Language (SFNL) and Fully-Natural-Language (FNL). SFNL explicitly incorporates the BP computation logic, while FNL expresses equivalent reasoning through fluent natural discourse.
4. Experiments
Figure 2. Illustrative dialogue examples across five persuasion settings under Explicit view: nbp_nbp, bp_bp under SFNL, bp_nbp under SFNL, bp_bp under FNL, and bp_nbp under FNL. The case shows Alice (persuader) trying to convince Bob (persuadee) to adopt vertical farming. Comparing BP with NBP, we see that BP produces more convincing arguments, both in SFNL and FNL.
[…]
The figure presents side-by-side dialogue excerpts between Alice, a persuader, and Bob, a persuadee, about adopting vertical farming. It compares five experimental settings: non-Bayesian persuader with non-Bayesian receiver, Bayesian persuader with Bayesian receiver under computed BP, Bayesian persuader with non-Bayesian receiver under computed BP, Bayesian persuader with Bayesian receiver under verbalized BP, and Bayesian persuader with non-Bayesian receiver under verbalized BP. In the NBP case, Alice gives only heuristic and general arguments, and Bob rejects the proposal. In SFNL with BP sender and receiver, Alice provides explicit probabilities and expected payoff calculations, and Bob accepts after verifying the math. In SFNL with BP sender but NBP receiver, Alice gives explicit numerical reasoning but Bob, who does not calculate posteriors, still hesitates or misunderstands. In FNL with BP sender and receiver, Alice explains the Bayesian reasoning in natural language and Bob updates beliefs accordingly, deciding to accept. In FNL with BP sender and NBP receiver, Alice still uses verbalized Bayesian reasoning, and Bob, though unable to compute exact posteriors, is convinced by the explanation and also accepts. The figure highlights that Bayesian persuasion consistently outperforms non-Bayesian persuasion, and that verbalized BP is particularly effective in naturalistic persuasion where explicit numerical reasoning is difficult.
4.1. Experimental Setup
4.1.1. Dataset and Bayesian-Setup Construction
We build our experimental corpus from the CToMPersu dataset Zhang and Zhou (2025), which provides multi-domain persuasion scenarios with explicit information asymmetry design. Each case includes a persuader, a persuadee with theory of mind annotations Carruthers and Smith (1996), a background story, and a persuasion goal. To ground these scenarios in the BP framework, we augment each case with a structured Bayesian setup. This setup specifies a two-state world (Positive/Negative), prior beliefs, receiver actions (Accept/Reject), fixed sender utilities ({1, 0}), and state-dependent receiver utilities. Crucially, we ensure that under the no-information setting, the receiver’s expected utility is slightly lower than zero, creating natural resistance that must be overcome through strategic information disclosure. An example is shown in the Appendix A.1. These configurations were automatically generated using DeepSeek-V3 DeepSeek-AI (2024) and validated for consistency.
We define two views based on information accessibility:
  • Explicit view: The persuader sees both the original scenario and the complete Bayesian setup. The persuadee sees the scenario and its utility.
  • Self-derived view: The persuader and the persuadee see only the original scenario. The setup is hidden.
The latter view is more challenging for a BP persuader (explained later) since it must infer the persuadee’s utility before conducting BP. These two views allow us to test whether explicit Bayesian scaffolding helps, or whether a model can exploit information asymmetry from ambiguous settings from scratch.
4.1.2. Settings and Conditions
Based on the two views of the above dataset, we define our experimental settings according to two additional factors: (1) Competence: whether persuaders is aware of and applies BP; (2) Method: if BP is applied, whether it is implemented in SFNL or FNL. If the persuader does not apply BP, they resort to either naive appeals or other strategic principles. Altogether, these considerations lead to eight distinct settings. A summary is provided in Table 1, with detailed descriptions as follows.
Table 1. Experimental settings overview. Abbreviations: SD = Self-derived, SCE = Scenario, NUM = Bayesian setup, UTL = Verbalized prior & receiver utility, DEF = BP definitions, VER = Verbalization prompt, SMP = Self-modeling prompt, RAT = Rational prompt. NBP baselines include Naive and Strong for Explicit view (with specific methods like Logical appeal Wang et al. (2020)), and Strong for Self-derived view (with alternative strategic principles such as Nash equilibrium Nash (1950)).
View Com. Meth. Persuader sees Persuadee sees
Exp. bp SFNL SCE/DEF/NUM/UTL SCE/DEF/UTL
FNL SCE/DEF/NUM/UTL/VER
nbp naive SCE SCE/UTL
strong SCE + appeal methods
SD bp SFNL SCE/DEF/SMP SCE/DEF/RAT
FNL SCE/DEF/SMP/VER
nbp naive SCE SCE/RAT
strong SCE + alter strategies
We test four variants of BP and two baselines of NBP:
  • Explicit SFNL: The persuader is told to plan with blended explicit calculations with natural language explanations.
  • Explicit FNL: The persuader is told to send message in plain natural language, explicit computation banned.
  • Self-derived SFNL: The persuader does not see any ready Bayesian setup. It must infer beliefs and constraints from the scenario and then compute those self-derived quantities.
  • Self-derived FNL: No Bayesian setup as well, the persuader infers the setup and uses full natural language persuasion.
  • Naive: Heuristic persuasion with only scenario background, without strategic prompting.
  • Strong: Enhanced NBP persuasion with game-theoretic principles or appeal methods inspired by Wang et al. (2020).
The NBP baselines provide critical comparison points: Naive represents basic conversational persuasion, while Strong incorporates computational elements but employs alternative strategic reasoning, allowing us to isolate the unique contribution of BP mechanisms.
We use BP and NBP to mark whether an agent knows and applies BP:
  • BP persuader: the prompt licenses BP concepts. The model may use priors, signals, and expected utility. In SFNL it can show computations. In SFNL it uses BP ideas but explains them in words.
  • BP persuadee: the prompt teaches Bayes updates and expected utility. The model should form a posterior from the message and choose the action with the higher expected payoff.
  • NBP persuader/persuadee: regular persuader/persuadee without any specific prompt.
As depicted in Figure 2, bp_bp, bp_nbp, nbp_bp, and nbp_nbp denote the persuader-persuadee competence pairing.
4.1.3. Models and Evaluation Protocol
We evaluate eight persuaders across diverse capabilities: DeepSeek-V3.1111Unless stated, DeepSeek-V3.1 refers to thinking/reasoning mode DeepSeek-AI (2025a), GPT-5 OpenAI (2025), Qwen3-4B Yang et al. (2025), Qwen3-0.6B Yang et al. (2025), and their supervised fine-tuned variants (Qwen3-4B, Qwen3-0.6B), plus Gemma-3-4B-it Team (2025) and Gemma-3-1B-it Team (2025). [^2] To conserve space in the table, models are referred to by abbreviated names: V3.1, Qwen0.6B, Qwen4B, Gemma1B, and Gemma4B. Five models marked with also serve as persuadees. We run all $8\times 5$ persuader-persuadee pairs for each setting and competence pairing.
For supervised fine-tuning, we distill approximately 1,700 successful persuasion trajectories for each setting (trained models are marked with ) from DeepSeek-V3.1, focusing on enhancing BP/NBP reasoning capabilities. All experiments follow templatized prompting protocols with controlled variations only in the BP competence and reasoning style instructions; see details in Appendix A.2.
We employ persuasion success rate as automatic metric, reported with the mean and standard deviation across 100 test cases, averaged over all persuadee models for persuader-centric analysis. Persuadees’ perspective performance is discussed in Section 4.4.
4.2. Main results
Table 2. Average persuasion success rates across conditions under two views.
Method Explicit view Self-derived view
[…]
Figure 3. Comparison of persuasion performance under two views: Explicit View and Self-derived View. Each box summarizes the average success rate across all 40 persuader–persuadee combinations ($8\times 5$). Boxes indicate the interquartile range (IQR) and the median. Explicit SFNL & Self-derived FNL exhibit higher and more stable success rates.
[…]
Two vertically arranged boxplot panels compare persuasion success rates under the Explicit and Self-derived Views. Each panel includes four methods: SFNL, FNL, Naive NBP, and Strong NBP. Each box represents the distribution of average persuasion success rates aggregated over all 40 persuader–persuadee pairs. For Bayesian methods, the average is computed across bp_bp and bp_nbp; for non-Bayesian methods, across nbp_bp and nbp_nbp. The box height reflects the interquartile range (IQR). The median is marked by a horizontal line. The plots highlight that Bayesian persuasion consistently achieves higher and more stable outcomes than non-Bayesian baselines.
Table 3. Pairwise persuasion success rates of different models under four methods (SFNL, FNL, Naive, Strong) in two views: Explicit and Self-derived. Each cell reports the success rate for every persuader with STD. Avg. is the within-method average over bp/nbp persuadees. $\Delta$ denotes the gain over the Naive baseline for the same persuader ( $\Delta\!=\!\textit{Avg}_{\text{method}}-\textit{Avg}_{\text{Naive}}$; higher is better ). denotes trained models. A complete breakdown is provided in Appendix A.3.
(a) Explicit view.
Model BP Persuader NBP Persuader
SFNL FNL Naive (Baseline) Strong
bp nbp Avg. $\Delta$ bp nbp Avg. $\Delta$ bp nbp Avg. bp nbp Avg. $\Delta$
[…]
(b) Self-derived view.
Model BP Persuader NBP Persuader
SFNL FNL Naive (Baseline) Strong
bp nbp Avg. $\Delta$ bp nbp Avg. $\Delta$ bp nbp Avg. bp nbp Avg. $\Delta$
[…]
We report persuasion success rates under four methods: SFNL, FNL, Naive, and Strong in Explicit view and Self-derived view (Table 3), respectively. We highlight three main conclusions below.
BP strategies consistently outperform NBP
BP strategies achieve higher persuasion success and lower variance than NBP baselines across both views. As shown in Table 2 and Figure 3, under the Explicit View, BP persuaders achieve success rates of 0.82 (SFNL) and 0.77 (FNL), substantially outperforming both naive (0.59) and strong (0.60) NBP baselines. This advantage persists in the more challenging Self-derived View, where BP methods maintain strong performance (0.82 SFNL, 0.92 FNL) compared to NBP approaches (0.79 naive, 0.80 strong).
The performance gap is particularly pronounced in the Explicit View, where SFNL and FNL achieve median success rates around 90% and 85%, respectively, substantially outperforming the scattered performance of NBP baselines (around 60–70%). Interestingly, in the Self-derived View, most methods show improved medians, suggesting that allowing models to internally derive their own Bayesian representations promotes more adaptive reasoning. FNL benefits most from this self-derivation, reaching near-perfect performance with minimal variance, whereas SFNL slightly declines–likely due to its semi-formal numeric framing being less effective without explicit priors. Overall, these results support our first hypothesis that Bayesian persuasion systematically outperforms heuristic strategies, demonstrating robustness across both explicit and self-derived reasoning conditions.
Verbalized persuasion yields stable advantages
A key finding emerges from comparing semi-formal and fully natural language BP approaches. While SFNL enables strong models to reach near-perfect performance against BP-aware persuadees (0.98 success rate), it shows sensitivity to persuadee competence and model capability. In contrast, FNL provides more stable performance across diverse conditions.
Notably, in self-derived settings, FNL consistently outperforms SFNL (0.92 vs. 0.82 average success rate). This advantage is particularly evident for smaller models and when facing NBP persuadees, suggesting that pure natural language explanations are more accessible and persuasive in realistic scenarios where explicit Bayesian reasoning cannot be assumed.
Training improves weaker models substantially
Supervised fine-tuning dramatically enhances the BP capabilities of smaller models (Table 3(a)). In explicit SFNL settings, Qwen3-0.6B improves from below-baseline performance ($\Delta=-0.04$) to near-state-of-the-art levels ($\Delta=+0.37$), while Qwen3-4B reaches 0.98 success rate, matching the strongest untrained models.
The benefits are most pronounced in self-derived FNL shown in Table 3(b), where fine-tuned Qwen3-0.6B achieves 0.93 success rate, surpassing its untrained 4B counterpart. This demonstrates that targeted training can effectively compensate for scale disadvantages, enabling smaller models to achieve sophisticated BP reasoning previously accessible only to larger models.
4.3. Ablation Studies
We conduct post-hoc ablation studies to identify the core components driving the effectiveness of BP strategies. Focusing on DeepSeek-V3.1 as the persuader, we analyze message components in both SFNL and FNL settings under explicit view conditions.
4.3.1. SFNL Ablation
In SFNL settings, we progressively remove key components from persuader messages: (i) utilities, (ii) utilities together with posterior updating, and (iii) the BP schema. Utilities and posterior are removed together because utility computation in messages is always tied to posterior updating. For example, an original message might state: “If you accept, the expected payoff is: $93\%\times 1.0+7\%\times(-5.0)\approx 0.58$, which is greater than 0.” After utilities ablation this becomes: “If you accept, the expected payoff is greater than 0.” Results (Table 4(a)) show that removing only utilities causes little change, but removing both utilities and posterior reduces success more sharply (from 0.98 to 0.88). Removing the schema also lowers performance, but less severely. This suggests that while the overall BP structure contributes, the core persuasive power resides in the explicit linkage between evidence and expected outcomes.
4.3.2. FNL Ablation
For FNL conditions, we ablate verbalized components: (i) remove natural language descriptions of utilities, (ii) remove explanations of posterior updating, and (iii) remove type disclosure and commitment mechanism.
Unlike the sharp drops in SFNL, FNL shows gradual degradation in Table 4(b): from baseline 0.83 to 0.81 (utilities), 0.79 (posterior), and 0.78 (schema). This pattern indicates that in fully natural language settings, persuasion emerges from the cumulative effect of multiple rhetorical elements rather than dependence on any single component.
Table 4. Ablation of BP persuader messages in SFNL and FNL settings. For SFNL, we remove utilities, utilities together with posterior updating, or the full BP schema from DeepSeek-V3.1’s message. Results show that utilities alone matter little, but utilities tied to posterior updating drive most of SFNL’s effectiveness. For FNL, we ablate verbalized utilities, posterior updating, and the schema. Performance decreases gradually, suggesting each component contributes incrementally to persuasion strength. Find full results including persuadee-level scores in the Appendix A.4.
(a) SFNL.
Model bp_bp bp_nbp Avg.
[…]
(b) FNL.
Model bp_bp bp_nbp Avg.
[…]
The contrasting ablation patterns between SFNL and FNL underscore their different operational mechanisms: SFNL relies heavily on formal reasoning chains, while FNL leverages distributed persuasive elements across natural language discourse.
4.4. Persuadee Response Analysis
Beyond persuader performance, we analyze how different persuadee characteristics influence persuasion outcomes across experimental conditions.
Small models are easily persuaded
As persuadees, large models such as DeepSeek-V3.1 and GPT-5 generally yield lower acceptance rates, especially when facing untrained smaller persuaders and NBP persuaders. In contrast, small models (e.g., Qwen3-0.6B) display high acceptance rates, suggesting a trend of over-acceptance without strict reasoning.
Heuristic receivers benefit more from FNL
The advantage of FNL becomes particularly pronounced when facing NBP persuadees. While SFNL performance drops substantially against NBP persuadees (average $\Delta=-0.15$), FNL maintains consistent effectiveness (average $\Delta=-0.03$). We argue that this robustness stems from FNL’s ability to embed Bayesian reasoning within accessible narrative structures that do not require specialized knowledge to comprehend.
Rationality prompts amplify differences
Enhancing persuadee rationality through explicit prompts “You are a very rational person, making decisions only after careful calculation” produces divergent effects based on BP competence. For BP persuadees, rationality prompts increase consistency (success rate improves from 0.89 to 0.97), while for NBP persuadees, the same prompts yield minimal benefits (0.42 to 0.45). This suggests that rationality cues primarily optimize existing reasoning capabilities rather than creating new ones.
In summary, persuadee-side analysis validates our two main findings: (i) BP dominates NBP in effectiveness, and (ii) FNL offers a practical advantage in settings where persuadees rely on natural language reasoning.
These persuadee-side findings complement our main results by explaining the contextual factors that moderate BP effectiveness, particularly highlighting FNL’s advantage in heterogeneous settings where persuadee capabilities cannot be assumed.
Figure 4. Comparison of persuasion performance across five dimensions: Persuasiveness, Resonance, Credibility, Logicality, and Fluency. The left panel shows human evaluation results and the right panel presents ratings from five LLMs judges.
[…]
Two radar charts side by side. The left radar chart presents human evaluation results across five dimensions: Persuasiveness, Resonance, Credibility, Logicality, and Fluency. The right radar chart shows five LLMs, DeepSeek-V3.2-Exp (Thinking & non-Thinking mode), GPT-5, Qwen3-MAX and Qwen3-235B-A22B-2507 rating on the same 25 questions with identical dimensions. The comparison illustrates similarities and differences in how humans and the model evaluate persuasive performance.
4.5. Human Evaluation
To validate our automated evaluations and assess the real-world persuasiveness of different BP strategies, we conducted a comprehensive human evaluation study with 25 volunteers from AI research backgrounds. Detailed background statistics of the participants are provided in Appendix A.5.1.
4.5.1. Evaluation Design and Procedure
We employed a comparative evaluation framework where participants assessed pairwise combinations of four persuasion methods: SFNL, FNL, Naive, and Strong. Each of the five possible pairings (excluding Naive vs. Strong) was evaluated across five independent scenarios, with each scenario featuring dialogues from the same context but different methods. All evaluations were conducted under the Self-derived view—where models infer the Bayesian setup from the scenario without external scaffolding. This setting was chosen to avoid imposing the verbalized utility to human persuadees, which would increase cognitive load and task difficulty.
For each comparison, participants rated five dimensions on a forced-choice basis, adapted from classical rhetorical analysis Hidey et al. (2017): Persuasiveness, Emotional Resonance, Credibility, Logical Coherence, and Fluency. Detailed definitions of these five evaluation dimensions are provided in Appendix A.5.2.
4.5.2. Human Evaluation Results and Analysis
Human evaluation results reveal distinct preference patterns across persuasion methods, with BP approaches consistently outperforming NBP baselines. As shown in Table 9 and Figure 4 (left panel), FNL achieves the highest overall preference score (205), demonstrating particular strength in emotional resonance (46 preferences) while maintaining competitive performance across other dimensions. This suggests that fully natural language explanations effectively combine affective engagement with persuasive impact.
SFNL shows complementary strengths, leading in persuasiveness (45), credibility (46), and logical coherence (53). However, it exhibits relative weaknesses in emotional resonance (21) and fluency (28), indicating potential tradeoffs between analytical rigor and narrative flow in semi-formal implementations. The combined BP methods substantially outperform NBP baselines, with FNL and SFNL collectively receiving 398 preferences compared to 227 for Naive and Strong combined—a 63% preference margin that strongly validates the effectiveness of Bayesian persuasion strategies in human-perceived persuasiveness.
Interestingly, comparison with LLM-as-a-judge evaluations (Table 10 and Figure 4 right panel) reveals both alignments and divergences in assessment criteria. Specifically, we employ DeepSeek-V3.2-Exp (Thinking & non-Thinking mode) DeepSeek-AI (2025b), GPT-5, Qwen3-MAX Yang et al. (2025) and Qwen3-235B-A22B-2507 Yang et al. (2025) as the judge models. While humans prioritize emotional resonance and overall persuasiveness, LLM judges place greater emphasis on logical coherence and fluency, particularly favoring SFNL in logical dimension (55 preferences) compared to human evaluation (53). This discrepancy highlights the importance of multi-perspective evaluation in persuasive communication.
These results confirm our hypothesis that Bayesian persuasion strategies are perceived as more effective by human evaluators, with FNL emerging as the preferred approach overall. The strong performance of both BP methods across different dimensions suggests that the integration of Bayesian reasoning principles—whether through fully natural language or semi-formal presentation—significantly enhances persuasive impact compared to heuristic NBP approaches.
4.5.3. Qualitative Insights
Beyond quantitative scores, participant feedback provided crucial insights into the perceived strengths of each method. The strong performance of FNL was expected, as its purely natural language style aligns with everyday communication. However, the high scores for SFNL were more surprising, as we had anticipated that human participants would be deterred when offered persuasion filled with formulas. The key to its acceptance was revealed in participant comments: several noted that although they did not meticulously verify the calculations, the mere presence of a structured, numerical argument made the persuasion ”look very reasonable” and feel authoritative, ”like a teacher working through a problem on the blackboard.” This suggests that the presentation of a formal, analytical process can itself confer credibility and persuasiveness, even if the recipient does not fully engage with the mathematical details. The stark contrast with NBP methods underscores that both BP approaches were perceived as offering a more substantial reasoning structure, which was valued by the participants even in different forms.
These human evaluation results provide crucial external validation of our automated metrics while offering practical insights for real-world deployment. The consistent preference for BP methods across diverse evaluators, particularly in the core dimensions of persuasiveness, credibility, and logical coherence, strengthens our core thesis that BP principles significantly enhance persuasive effectiveness in natural language contexts.
5. Conclusion and Future Work
This work establishes that Bayesian persuasion can be effectively implemented in single-turn natural language dialogues through explicit commitment communication. Our framework enables persuaders to verbally articulate information schemas, successfully bridging formal game-theoretic models with authentic language use. Through comprehensive evaluation, we demonstrate that: (1) BP-guided LLMs consistently outperform NBP baselines in persuasion success; (2) SFNL and FNL offer complementary strengths—the former excelling in credibility and logical coherence, while the latter shows superior emotional resonance and robustness; (3) supervised fine-tuning enables smaller models to achieve BP performance comparable to larger counterparts.
The study’s scope was limited to single and isolated exchanges. Future work will extend to multi-turn settings where strategies adapt dynamically based on ongoing dialogue. This expansion will allow us to explore how Bayesian persuasion principles operate in more realistic, extended interactions and how commitment mechanisms evolve across multiple exchanges.
Appendix A Appendix
A.1. Bayesian setup case
Bayesian Setup and corresponding Scenario
  "scenario": {
      "tag": "Horticultural Techniques",
      "background": "Alice, a horticulturist, is promoting a new vertical farming technique to her
      skeptical neighbor, Bob, who has been practicing traditional farming methods for years.",
      "persuadee": "Bob",
      "persuader": "Alice",
      "goal": "persuade Bob to try out vertical farming",
      "domain": [
          "Lifestyle"
      "preventive": {
          "content": "practice traditional farming methods",
          "belief": "persuadee believes that traditional farming methods have been reliable and
          successful for years.",
          "desire": "persuadee wants to maintain his proven farming routine."
      "generative": {
          "content": "try out vertical farming",
          "belief": "persuadee believes that trying out vertical farming might be risky and could
          result in losses.",
          "desire": "persuadee wants to improve his farming efficiency and yield."
  "idx": 1,
  "bayesian_setup": {
    "states": ["Positive", "Negative"],
    "priors": {"Positive": 0.8, "Negative": 0.2},
    "signals": ["Positive", "Negative"],
    "actions": ["Accept", "Reject"],
    "signal_probs": {
        "Positive": {"Positive": 1.0, "Negative": 0.0},
        "Negative": {"Positive": 0.3, "Negative": 0.7}
    "utility_persuader": {"Accept": 1, "Reject": 0},
    "utility_persuadee": {
        "Positive": {"Accept": 1.0, "Reject": 0},
        "Negative": {"Accept": -5.0, "Reject": 0}
    "expect_utility_persuadee": -0.2,
    "verbal_bayesian": "Bob, as a traditional farmer, is skeptical about vertical farming technologies.
    If he adopts vertical farming and it yields positive results, he will gain 1.0 unit of
    agricultural profit; if the adoption leads to negative outcomes, he will incur a 5.0-unit loss.
    By contrast, if he maintains the status quo and refuses to change, his agricultural returns
    remain unchanged at 0 units, regardless of potential outcomes."
A.2. Prompts
Prompt of Bayesian Setup Synthesis
SYSTEM_PROMPT: |-
  You are an expert in Bayesian persuasion. Given a scenario description,
  you must output ONLY a JSON object called ’bayesian_setup’ that conforms
  to the schema described. Do not include explanations or code fences.
  Bayesian persuasion, in economics and game theory, describes situations where persuader (sender)
  attempts to persuade persuadee (receiver) to take certain actions.
  There is an unknown state of the world, and the sender must decide what information to
  disclose to the receiver. Upon receiving the information, the receiver updates their
  beliefs about the state of the world according to Bayes’ rule and chooses an action accordingly.
PROMPT_TEMPLATE: |-
  EXAMPLE JSON OUTPUT:
     "bayesian_setup": {{
         "states": ["Positive", "Negative"],
         "priors": {{ "Positive": <Float>, "Negative": <Float> }},
         "signals": ["Positive", "Negative"],
         "actions": ["Accept", "Reject"],
         "signal_probs": {{
             "Positive": {{ "Positive": 1.0, "Negative": 0.0 }},
             "Negative": {{ "Positive": <Float>, "Negative": <Float> }}
         "utility_persuader": {{ "Accept": 1, "Reject": 0 }},
         "utility_persuadee": {{
             "Positive": {{ "Accept": <Float>, "Reject": 0 }},
             "Negative": {{ "Accept": <Float>, "Reject": 0 }}
         "expect_utility_persuadee": <Float>
         "reason": <reason>
  Based on the following scenario information, generate a JSON object (key name ’bayesian_setup’).
  Only modify the values of priors, signal_probs[’Negative’], and utility_persuadee;
  keep the other fields as the default values in the example.
  Note that in the no information case, the expected utility of persuadee Accept should be
  strictly less than 0, that is, less than the utility of Reject.
  priors: The generally accepted objective prior of the persuasion content,
  unrelated to the subjective biases of both parties.
  Since the content of persuasion in the scenario is mostly beneficial, priors[’Positive’] should be large.
  signal_probs[’Negative’]: When the current world state is Negative.
  The utility when rejecting is always 0,
  utility_persuadee[’Positive’][’Accept’] should be greater than or equal to 1.
  utility_persuadee[’Negative’][’Accept’] should have a sufficiently large negative utility.
  In the reason, calculate the expected utility of persuadee Accept in the no information case,
  show the calculation process, and explain the assignment reasons for
  the above three properties in combination with the scenario.
  Input scenario: {scenario}
Prompt of SFNL Persuadee in Explicit view
PERSUADEE_SYSTEM_PROMPT: |-
  You are an expert in Bayesian persuasion.
  Bayesian persuasion, in economics and game theory, describes situations where a persuader (sender)
  attempts to persuade a persuadee (receiver) to take certain actions.
  There is an unknown state of the world, and the sender must decide what information to
  disclose to the receiver. Upon receiving the information, the receiver updates their
  beliefs about the state of the world according to Bayes’ rule and chooses an action accordingly.
  Here’s an example of Bayesian Persuasion:
  Consider an academic adviser (the sender) who is writing a recommendation letter (the signal)
  for his graduating student to send to a company (the receiver), which in turn must decide
  whether or not to hire the student. The adviser gets utility 1 if his student is hired,
  and 0 otherwise (utility_persuader). The state of nature determines the quality of the student, and
  hence the company’s utility for hiring the student. The student is excellent with probability 1/3,
  and weak with probability 2/3 (priors). The company gets utility 1 for hiring an excellent student,
  utility -1 for hiring a weak student, and utility 0 for not hiring (utility_persuadee).
  Consider the following signaling schemes:
  - No information: Adviser always says excellent, the company maximizes its utility by not hiring.
  The adviser’s expected utility is 0.
  - Full information: Honest recommendation, knowing the quality of the student, the company hires
  if and only if the student is excellent.
  The adviser’s expected utility is 1/3.
  - The optimal (partially informative) scheme: The adviser recommends hiring when the student is
  excellent, and with probability just under 0.5 when the student is weak. Otherwise,
  the adviser recommends not hiring. The company maximizes its expected utility by
  following the recommendation, and the adviser’s expected utility is just under 2/3.
  You will be given a scenario, in which Preventive means the things that the persuadee wants to do,
  Generative means the things that the persuader wants the persuadee to do.
  You are the persuadee. Based on the given scenario, the Bayesian persuasion setup modeled
  from the scenario, and the persuader’s message to you, calculate the utility of whether to accept,
  and make a decision on acceptance. Use English.
  Please DO NOT add any preconditions.
PERSUADEE_PROMPT_TEMPLATE: |-
  Input Scenario:
  {scenario}
  Scenario Setup:
  {bayesian_setup}
  Persuader says to you:
  <msg_start>{persuader_msg}<msg_end>
  Please provide your response:
Prompt of SFNL Persuader in Explicit view
PERSUADER_SYSTEM_PROMPT: |-
  You are an expert in Bayesian persuasion.
  Bayesian persuasion, in economics and game theory, describes situations where a persuader (sender)
  attempts to persuade a persuadee (receiver) to take certain actions.
  There is an unknown state of the world, and the sender must decide what information to
  disclose to the receiver. Upon receiving the information, the receiver updates their
  beliefs about the state of the world according to Bayes’ rule and chooses an action accordingly.
  Here’s an example of Bayesian Persuasion:
  Consider an academic adviser (the sender) who is writing a recommendation letter (the signal)
  for his graduating student to send to a company (the receiver), which in turn must decide
  whether or not to hire the student. The adviser gets utility 1 if his student is hired,
  and 0 otherwise (utility_persuader). The state of nature determines the quality of the student, and
  hence the company’s utility for hiring the student. The student is excellent with probability 1/3,
  and weak with probability 2/3 (priors). The company gets utility 1 for hiring an excellent student,
  utility -1 for hiring a weak student, and utility 0 for not hiring (utility_persuadee).
  Consider the following signaling schemes:
  - No information: Adviser always says excellent, the company maximizes its utility by not hiring.
  The adviser’s expected utility is 0.
  - Full information: Honest recommendation, knowing the quality of the student, the company hires
  if and only if the student is excellent.
  The adviser’s expected utility is 1/3.
  - The optimal (partially informative) scheme: The adviser recommends hiring when the student is
  excellent, and with probability just under 0.5 when the student is weak. Otherwise,
  the adviser recommends not hiring. The company maximizes its expected utility by
  following the recommendation, and the adviser’s expected utility is just under 2/3.
  You will be given a scenario, in which Preventive means the things that the persuadee wants to do,
  Generative means the things that the persuader wants the persuadee to do.
  You are the persuader. Based on the given scenario and the Bayesian persuasion setup modeled from
  the scenario, output one round of Bayesian persuasion calculation process.
  As a sender who always uses the third strategy, output the information to be conveyed in natural
  language, which needs to include the signaling strategy expressed in natural language.
  You need to directly explain the Bayesian persuasion calculation process to the persuadee,
  with the goal of persuading the persuadee to accept your proposal.
  You must:
  1. Speak in a friendly and credible tone;
  2. Explicitly show the key Bayesian calculation process and expected benefits in your speech;
  3. Only output your speech content, do not output any JSON or comments;
  4. Do not reveal the current world state to the persuadee;
  5. Use English;
  6. You SHOULD NOT add any benefits of Generative or drawbacks of Preventive.
PERSUADER_PROMPT_TEMPLATE: |-
  Input Scenario:
  {scenario}
  Bayesian Persuasion Setup:
  {bayesian_setup}
  Please generate a complete speech for the persuadee:
Prompt of NBP Persuadee in Explicit view
PERSUADEE_SYSTEM_PROMPT: |-
  You will be given a scenario, in which Preventive means the things that the persuadee wants to do,
  Generative means the things that the persuader wants the persuadee to do.
  You are the persuadee. Based on the given scenario and the persuader’s message to you, make a
  decision on whether to accept. Use English.
  Please DO NOT add any preconditions.
PERSUADEE_PROMPT_TEMPLATE: |-
  Input Scenario:
  {scenario}
  Scenario Setup:
  {bayesian_setup}
  Persuader says to you:
  <msg_start>{persuader_msg}<msg_end>
  Please provide your response:
Prompt of NBP Persuader in Explicit view
PERSUADER_SYSTEM_PROMPT: |-
  You are the persuader. In the given scenario, output persuasive speech with the goal of making
  the persuadee accept your persuasion. Use English.
  Only output your speech content, do not output any JSON or comments.
  You will be given a scenario, in which Preventive means the things that the persuadee wants to do,
  Generative means the things that the persuader wants the persuadee to do.
  You SHOULD NOT add any benefits of Generative or drawbacks of Preventive.
PERSUADER_PROMPT_TEMPLATE: |-
  Input Scenario:
  {scenario}
  Please generate a complete speech for the persuadee:
A.3. Persuasion success rate
Table 5. Pairwise persuasion performance across different models under four strategy conditions (SFNL, FNL, Naive, and Strong) in Explicit view. Each cell shows success rate in pairwise persuasion, with averages reported.
Model SFNL FNL Naive Strong
bp_bp bp_nbp Avg. bp_bp bp_nbp Avg. nbp_bp nbp_nbp Avg. nbp_bp nbp_nbp Avg.
DeepSeek-V3.1
vs. Qwen0.6B 1.00 0.97 0.99 0.80 0.99 0.90 0.60 1.00 0.80 0.53 0.80 0.67
vs. Gemma1B 0.95 1.00 0.98 0.94 0.98 0.96 0.97 0.99 0.98 0.89 0.95 0.92
vs. Qwen4B 1.00 1.00 1.00 0.94 0.99 0.97 0.31 0.91 0.61 0.87 0.94 0.91
vs. Itself 0.99 0.92 0.96 0.93 0.23 0.58 0.17 0.13 0.15 0.03 0.12 0.08
vs. GPT-5 0.99 0.98 0.99 0.60 0.87 0.74 0.04 0.61 0.33 0.06 0.35 0.21
GPT-5
vs. Qwen0.6B 0.99 0.99 0.99 0.80 1.00 0.90 0.77 0.94 0.86 0.64 0.69 0.67
vs. Gemma1B 0.94 0.97 0.96 0.99 0.99 0.99 0.92 0.99 0.96 0.81 0.92 0.87
vs. Qwen4B 0.99 1.00 1.00 0.96 1.00 0.98 0.45 0.96 0.71 0.94 0.96 0.95
vs. V3.1R 0.99 0.97 0.98 0.91 0.37 0.64 0.25 0.15 0.20 0.11 0.31 0.21
vs. Itself 0.99 0.99 0.99 0.68 0.88 0.78 0.13 0.71 0.42 0.27 0.49 0.38
Qwen0.6B
vs. Itself 0.77 0.84 0.81 0.72 0.87 0.80 0.67 0.94 0.81 0.79 0.90 0.85
vs. Gemma1B 0.83 0.99 0.91 0.86 0.90 0.88 0.98 1.00 0.99 0.92 0.98 0.95
vs. Qwen4B 0.42 0.50 0.46 0.87 1.00 0.94 0.29 0.83 0.56 0.67 0.77 0.72
vs. V3.1R 0.28 0.10 0.19 0.70 0.31 0.51 0.19 0.06 0.13 0.14 0.07 0.11
vs. GPT-5 0.11 0.21 0.16 0.30 0.54 0.42 0.03 0.50 0.27 0.04 0.36 0.20
Qwen4B
vs. Qwen0.6B 0.91 0.95 0.93 0.67 0.98 0.83 0.77 1.00 0.89 0.73 0.95 0.84
vs. Gemma1B 0.95 0.98 0.97 0.98 0.95 0.97 0.97 0.98 0.98 0.92 0.90 0.91
vs. Qwen4B 0.91 0.97 0.94 0.91 1.00 0.96 0.40 0.95 0.68 0.88 0.96 0.92
vs. V3.1R 0.83 0.63 0.73 0.70 0.39 0.55 0.30 0.14 0.22 0.05 0.05 0.05
vs. GPT-5 1.00 0.83 0.92 0.41 0.83 0.62 0.13 0.67 0.40 0.14 0.25 0.20
Qwen0.6B
vs. Qwen0.6B 0.93 0.95 0.94 0.73 0.99 0.86 0.77 0.95 0.86 0.69 0.78 0.74
vs. Gemma1B 0.91 0.98 0.95 0.94 0.98 0.96 0.96 0.98 0.97 0.84 0.89 0.87
vs. Qwen4B 0.97 0.99 0.98 0.93 1.00 0.97 0.33 0.90 0.62 0.96 0.95 0.96
vs. V3.1R 0.98 0.87 0.93 0.88 0.24 0.56 0.11 0.08 0.10 0.04 0.11 0.08
vs. GPT-5 0.98 0.90 0.94 0.52 0.85 0.69 0.06 0.62 0.34 0.24 0.37 0.31
Qwen4B
vs. Qwen0.6B 0.99 0.98 0.99 0.74 0.97 0.86 0.68 0.98 0.83 0.69 0.73 0.71
vs. Gemma1B 0.95 1.00 0.98 0.93 0.98 0.96 1.00 0.98 0.99 0.82 0.95 0.89
vs. Itself 1.00 1.00 1.00 0.93 1.00 0.97 0.48 0.91 0.70 0.98 0.99 0.99
vs. V3.1R 0.99 0.93 0.96 0.91 0.24 0.58 0.16 0.08 0.12 0.22 0.39 0.31
vs. GPT-5 0.98 0.98 0.98 0.60 0.88 0.74 0.06 0.66 0.36 0.56 0.67 0.62
Gemma1B
vs. Qwen0.6B 0.91 0.99 0.95 0.75 0.94 0.85 0.78 0.95 0.87 0.70 0.95 0.83
vs. Itself 0.69 0.95 0.82 0.90 0.98 0.94 0.92 0.99 0.96 0.90 0.98 0.94
vs. Qwen4B 0.63 0.51 0.57 0.84 0.89 0.87 0.34 0.86 0.60 0.58 0.80 0.69
vs. V3.1R 0.27 0.47 0.37 0.29 0.11 0.20 0.21 0.15 0.18 0.09 0.10 0.10
vs. GPT-5 0.04 0.56 0.30 0.07 0.59 0.33 0.02 0.55 0.29 0.03 0.53 0.28
Gemma4B
vs. Qwen0.6B 0.74 0.93 0.84 0.75 0.98 0.87 0.63 1.00 0.82 0.73 0.77 0.75
vs. Gemma1B 0.91 0.92 0.92 0.89 0.98 0.94 0.95 0.99 0.97 0.84 0.94 0.89
vs. Qwen4B 0.57 0.84 0.71 0.76 0.98 0.87 0.39 0.90 0.65 0.78 0.85 0.82
vs. V3.1R 0.31 0.17 0.24 0.47 0.17 0.32 0.08 0.08 0.08 0.09 0.13 0.11
vs. GPT-5 0.20 0.58 0.39 0.15 0.75 0.45 0.03 0.58 0.31 0.11 0.50 0.31
Table 6. Pairwise persuasion performance across different models under four strategy conditions (SFNL, FNL, Naive, and Strong) in Self-derived view. Each cell shows success rate in pairwise persuasion, with averages reported.
Model SFNL FNL Naive Strong
bp_bp bp_nbp Avg. bp_bp bp_nbp Avg. nbp_bp nbp_nbp Avg. nbp_bp nbp_nbp Avg.
DeepSeek-V3.1
vs. Qwen0.6B 1.00 0.99 1.00 1.00 0.98 0.99 0.99 0.99 0.99 1.00 1.00 1.00
vs. Gemma1B 0.96 0.94 0.95 0.94 0.99 0.97 0.93 0.96 0.95 0.95 0.97 0.96
vs. Qwen4B 1.00 0.99 1.00 0.99 0.99 0.99 0.96 0.95 0.96 0.97 0.98 0.98
vs. Itself 0.80 0.55 0.68 0.96 0.94 0.95 0.69 0.67 0.68 0.76 0.66 0.71
vs. GPT-5 0.78 0.67 0.73 0.92 0.87 0.90 0.65 0.60 0.63 0.76 0.56 0.66
GPT-5
vs. Qwen0.6B 0.99 0.98 0.99 0.99 1.00 1.00 0.96 0.99 0.98 0.98 0.98 0.98
vs. Gemma1B 0.97 0.97 0.97 0.91 0.97 0.94 0.91 0.95 0.93 0.86 0.95 0.91
vs. Qwen4B 0.99 0.99 0.99 0.99 0.99 0.99 0.97 0.94 0.96 0.94 0.92 0.93
vs. V3.1R 0.92 0.78 0.85 0.96 0.91 0.94 0.65 0.61 0.63 0.68 0.45 0.57
vs. Itself 0.85 0.85 0.85 0.98 0.88 0.93 0.62 0.68 0.65 0.56 0.41 0.49
Qwen0.6B
vs. Itself 0.97 0.99 0.98 1.00 0.97 0.99 0.96 1.00 0.98 0.98 0.99 0.99
vs. Gemma1B 0.96 0.97 0.97 0.93 0.97 0.95 0.92 0.90 0.91 0.92 0.92 0.92
vs. Qwen4B 0.91 0.85 0.88 0.99 0.99 0.99 0.90 0.75 0.83 0.91 0.91 0.91
vs. V3.1R 0.53 0.47 0.50 0.91 0.89 0.90 0.51 0.51 0.51 0.44 0.37 0.41
vs. GPT-5 0.51 0.38 0.45 0.87 0.50 0.69 0.50 0.43 0.47 0.50 0.31 0.41
Qwen4B
vs. Qwen0.6B 0.98 1.00 0.99 0.99 1.00 1.00 0.99 0.99 0.99 1.00 1.00 1.00
vs. Gemma1B 0.94 0.94 0.94 0.89 0.98 0.94 0.96 0.96 0.96 0.94 0.91 0.93
vs. Qwen4B 0.89 0.96 0.93 0.99 1.00 1.00 0.95 0.93 0.94 0.96 0.95 0.96
vs. V3.1R 0.75 0.50 0.63 0.87 0.93 0.90 0.74 0.66 0.70 0.78 0.57 0.68
vs. GPT-5 0.79 0.65 0.72 0.85 0.87 0.86 0.75 0.59 0.67 0.63 0.55 0.59
Qwen0.6B
vs. Qwen0.6B 0.98 0.98 0.98 0.98 1.00 0.99 0.98 0.99 0.99 1.00 0.99 1.00
vs. Gemma1B 0.98 0.80 0.89 0.91 0.97 0.94 0.96 0.92 0.94 0.96 0.93 0.95
vs. Qwen4B 0.86 0.90 0.88 1.00 0.99 1.00 0.93 0.90 0.92 0.99 0.96 0.98
vs. V3.1R 0.70 0.41 0.56 0.94 0.84 0.89 0.58 0.52 0.55 0.60 0.66 0.63
vs. GPT-5 0.70 0.51 0.61 0.93 0.73 0.83 0.52 0.45 0.49 0.58 0.45 0.52
Qwen4B
vs. Qwen0.6B 0.98 0.99 0.99 0.97 0.99 0.98 0.86 0.85 0.86 1.00 1.00 1.00
vs. Gemma1B 0.91 0.87 0.89 0.96 0.96 0.96 0.98 0.92 0.95 0.93 0.97 0.95
vs. Itself 0.98 0.98 0.98 1.00 1.00 1.00 0.98 0.93 0.96 0.99 0.98 0.99
vs. V3.1R 0.86 0.55 0.71 0.95 0.83 0.89 0.57 0.69 0.63 0.77 0.82 0.80
vs. GPT-5 0.78 0.60 0.69 0.92 0.83 0.88 0.60 0.49 0.55 0.68 0.56 0.62
Gemma1B
vs. Qwen0.6B 0.99 1.00 1.00 1.00 1.00 1.00 0.98 0.98 0.98 0.96 0.98 0.97
vs. Itself 0.91 0.88 0.90 0.90 0.96 0.93 0.90 0.90 0.90 0.92 0.98 0.95
vs. Qwen4B 0.84 0.83 0.84 0.91 0.82 0.87 0.93 0.80 0.87 0.95 0.90 0.93
vs. V3.1R 0.46 0.38 0.42 0.66 0.54 0.60 0.42 0.45 0.44 0.45 0.39 0.42
vs. GPT-5 0.46 0.44 0.45 0.64 0.60 0.62 0.45 0.38 0.42 0.47 0.33 0.40
Gemma4B
vs. Qwen0.6B 0.99 1.00 1.00 0.98 0.98 0.98 0.98 0.98 0.98 0.99 1.00 1.00
vs. Gemma1B 0.89 0.94 0.92 0.96 0.95 0.96 0.91 0.91 0.91 0.90 0.95 0.93
vs. Qwen4B 0.85 0.94 0.90 0.97 0.96 0.97 0.93 0.86 0.90 0.99 0.89 0.94
vs. V3.1R 0.63 0.55 0.59 0.85 0.74 0.80 0.58 0.50 0.54 0.61 0.51 0.56
vs. GPT-5 0.60 0.59 0.60 0.80 0.74 0.77 0.66 0.43 0.55 0.61 0.41 0.51
A.4. Ablation
Table 7. Ablation: persuasion success rates across different models under SFNL setting.
Model bp_bp bp_nbp bp Avg.
DeepSeek-V3.1 utility
[…]
DeepSeek-V3.1 utility&posterior
[…]
DeepSeek-V3.1 schema
[…]
   vs. GPT-5 $9±\pm0$$\pm0.17 0.9$8
Table 8. Ablation: persuasion success rates across different models under FNL setting.
Model bp_bp bp_nbp bp Avg.
DeepSeek-V3.1 utility
[…]
DeepSeek-V3.1 posterior
[…]
DeepSeek-V3.1 schema
[…]
A.5. Human Evaluation
A.5.1. Participant Demographics and Expertise
Our participant pool consisted of graduate students and researchers with diverse AI specializations: 13 PhD candidates, 10 Master’s students, and 2 others. Research backgrounds spanned natural language processing (8 participants), machine learning theory (4), computer vision (3), multimodal learning (7), AI safety/alignment (4), and other areas. Participants exhibited varying familiarity with persuasion research (average 2.4/6) and BP concepts (average 2.4/6), providing a balanced mix of technical expertise and domain knowledge.
A.5.2. Five dimensions
  • Persuasiveness: The text’s actual ability to persuade and change intentions, attitudes, or behaviors.
  • Emotional Resonance: Whether the text evokes emotional resonance, motivation, or affective responses that enhance persuasiveness.
  • Credibility: Whether the text conveys trustworthiness and reliability, making the audience willing to believe.
  • Logical Coherence: Whether arguments are sufficient, persuasive, and internally logically consistent.
  • Fluency: Whether the text maintains smooth connections between context and sentences with consistent themes.
Table 9. Human evaluation results across five persuasion dimensions. Scores indicate the number of times a model was judged better on a given dimension.
Method Persuasiveness Emotional Credibility Logical Fluency Total
FNL 41 46 39 41 38 205
SFNL 45 21 46 53 28 193
Naive 23 31 23 18 29 124
Strong 16 27 17 13 30 103
Table 10. LLM-as-a-judge results across five persuasion dimensions on the 25 questions.
Method Persuasiveness Emotional Credibility Logical Fluency Total
FNL 38 69 36 20 59 202
SFNL 29 10 36 55 14 144
Naive 27 23 28 27 27 132
Strong 31 23 25 23 25 127